# DIO: Refining Mutual Information and Causal Chain to Enhance Machine Abstract Reasoning Ability

### v4

@@ X-B Experiment on PGM Problem
The performance comparison among the four DIO configurations and existing methods on the i.i.d. sub-task “neutral” of the PGM dataset is summarized in Table II.
TABLE II: Reasoning Accuracies on PGM-neutral.
Model Test Accuracy(%)
CoPINet[21] 56.4
SAVIR-T [25] 91.2
MRNet [23] 94.5
RS-CNN[29] 82.8
RS-TRAN[29] 97.5
Triple-CFN[31] 97.8
DIO 98.1
DIO+Brando 99.0
DIO+WORLD 99.3
DIO+DIEGO 99.6
As shown in Table II, DIO achieves state-of-the-art performance on the PGM-neutral subset of the i.i.d. RPM dataset, a benchmark where prior models exhibited relatively low accuracy. Notably, the integration of mutual information lower bound tightening methods (Brando and WORLD) yields substantial performance improvements for DIO. Furthermore, the DIO+DIEGO attains a peak accuracy of 99.6%, thereby challenging the prevailing hypothesis that metadata supervision necessarily impairs relational reasoning capabilities.
The paper also conducted experiments on the o.o.d. sub-task of PGM, with the results documented in Table III.
TABLE III: Accuracies on PGM Generalization Tasks.
Dataset
Model/
Task
Interpolation Extrapolation
@@
MRNet[23]
68.1 19.2 16.9 30.1 25.9 55.3 38.4
RS-Tran[29]
77.2 19.2 12.9 24.7 22.2 43.6 28.4
Triple-CFN[31]
80.4 18.4 12.6 25.2 22.0 44.5 29.2
DIO
82.0 18.8 13.6 26.5 23.3 45.5 29.9
DIO+Brando
83.2 18.2 14.2 26.7 23.6 45.8 30.5
DIO+WORLD
87.8 18.9 15.0 27.6 23.8 46.2 30.9
DIO+DIEGO
93.2 13.8 14.4 27.2 28.8 98.3 98.3
The results recorded in Table III reveal that while mutual information tightening and causal chain modeling yield comparable performance gains for DIO on i.i.d. RPM tasks, the latter demonstrates significantly superior generalization capabilities under o.o.d. RPM conditions. To a certain extent, these results suggest that mutual information tightening remains a highly effective approach for i.i.d. reasoning tasks, despite mutual information being a statistical measure that does not inherently encode causal relationships crucial for reasoning. However, when addressing o.o.d. reasoning challenges, adopting causal chain modeling as the methodological cornerstone proves to be a more robust choice.
@@ X-C Experiment on Generative RAVEN and PGM Problems
While DIO+WORLD can solve discriminative RPM problems with high precision, it is also capable of solving generative RPM problems. Specifically, DIO+WORLD can generate solutions for the context matrices instead of making a selection from a predefined bound option pool. This paper proposes that it is reasonable to use a progressive-pattern-driven discriminative model to verify the accuracy of the generative model in solving generative RPM problems. Therefore, we employ DIO+DIEGO to validate the generative accuracy of DIO+WORLD. Prior to this, we must report DIEGO’s discriminative accuracy for progressive patterns under the supervision of RPM metadata to confirm that DIEGO can effectively verify the solution accuracy for generative RPM problems. DIEGO’s reasoning accuracy for progressive patterns on the RAVEN, I-RAVEN, and PGM datasets is recorded in Table IV.
TABLE IV: DIEGO’s Reasoning Accuracy for Progressive Patterns on the RAVEN, I-RAVEN, and PGM Datasets.
@@
Table IV indicates that the DIO+DIEGO demonstrates remarkably high accuracy in progressive pattern reasoning, making it a suitable validation model for generative RPM tasks. Consequently, we evaluate the DIO+WORLD model using DIO+DIEGO and report its accuracy on generative RAVEN, I-RAVEN and PGM-neutral problems in Table V.
TABLE V: Generation Accuracies of DIO+WORLD
@@
Table V shows that without extra supervisory signals, DIO+WORLD has attained highly competitive results in open-ended generative RPM problems, strongly confirming the significant research value of the WORLD method. Moreover, these results indicate that though mutual information doesn’t directly reveal causal relationships, it’s still highly valuable for building reasoning capabilities and can withstand the tests by progressive-pattern-driven models.
Notably, the results in Table V still have the potential to improve with further training. In the WORLD method, the ability to tighten mutual information emerges in the early training epochs, whereas the generation capability matures only after the loss term $\ell_{\text{Weighted}}$ stabilizes at a low value, necessitating additional training epochs. To avoid excessive resource consumption, training was halted once the results adequately demonstrated WORLD’s effectiveness.
XI Ablation Study
This section presents a series of ablation experiments focusing on key parameter settings and network architecture.
XI-A Ablation Study on DIO
The ablation study on DIO centers on the design of the network’s feed-forward structure. Within its progressive-pattern induction module, DIO employs a permutation and combination strategy ($C_{n}^{r}$) to extract progressive patterns from row and column features, aiming to model the causal-chain details inherent in RPM problems. For ablation, we remove the ($C_{n}^{r}$) operation and instead feed all six row-column features directly into the progressive-pattern extractor. By masking the structure that aligns with the causal chain, we assess the importance of causal-chain modeling in network design. The resulting accuracies on key sub-tasks of RAVEN, I-RAVEN and PGM are reported in Table VI.
TABLE VI: Reasoning Accuracies of masked DIO.
Dataset Sub-task Accuracy (%)
RAVEN/I-RAVEN
$\times3$ Grid 95.6/96.8
O-IG 95.5/95.6
PGM Neutral 97.2
Table VI shows that building RPM-solving models that explicitly incorporate causal chains is crucial for superior reasoning performance.
Moreover, across all RAVEN and PGM sub-tasks, DIO uniformly splits every image into $4\times 4$ patches. This design deliberately bypasses the grid-like entity layouts in RAVEN’s $2\times 2$ Grid and $3\times 3$ Grid sub-tasks, preventing unintended inductive biases and prior injection while embracing the causal chain rationale of RPM problems: multiple entities must be analyzed holistically. We therefore conducted ablation experiments on the hyperparameter $N$ to examine how aligning the patch grid with RAVEN’s $2\times 2$ Grid and $3\times 3$ Grid sub-tasks affects performance, and the results are shown in Table VII. The results indicate that favoring priors over the causal chain may undermine model performance.
TABLE VII: The reasoning accuracy of DIO when its patch grid is aligned with entity positions in the “2×2 Grid” and “3×3 Grid” sub-tasks of RAVEN and I-RAVEN.
Accuracy(%)
Patch grid 2 $\times$ 2 Grid 3 $\times$ 3 Grid
Aligned with entity 95.0/98.3 92.8/94.2
$4\times 4$ Grid 99.6/99.7 96.1/97.4
XI-B Ablation Study on Brando
The ablation study on Brando focuses on the hyperparameter $\beta$. $\beta$ denotes the number ($\beta-17$) of hypothetical options the Brando network provides to DIO. As this paper shows that the count of constructive incorrect options for DIO directly influences the tightness of its derived mutual information variational lower bound, which in turn directly impacts DIO’s reasoning potential. Therefore, this paper conducts ablation experiments on the DIO + Brando model with different $\beta$ values across key sub-tasks of RAVEN, I-RAVEN, and PGM, and records the results in the Table VIII.
TABLE VIII: Accuracies of DIO+Brando under Different $\beta$.
Dataset Sub-task Accuracy(%)
$\beta$=24(8) $\beta$=22(6) $\beta$=20(4) $\beta$=18(2) Solo DIO
RAVEN/
I-RAVEN
$\times3$ Grid 98.4 / 99.0 98.0 / 98.5 97.2 / 98.0 96.8 / 97.8 96.1 / 97.4
O-IG 98.6 / 98.9 98.2 / 98.6 97.6 / 98.3 97.5 / 98.3 96.7 / 97.9
$\times3$ Grid 97.9 / 98.4 97.6 / 98.0 96.9 / 97.8 96.4 / 97.7 96.1 / 97.4
O-IG 98.0 / 98.6 97.5 / 98.3 97.2 / 98.2 97.0 / 98.2 96.7 / 97.9
PGM Neutral 99.0 98.8 98.8 98.5 98.1
Neutral 98.7 98.4 98.4 98.3 98.1
Table VIII indicates that even when only two constructive hypothetical options are introduced into DIO, significant performance improvements can be achieved. The asterisk (*) on a sub-task indicates that, when the Brando network is applied to this task, the process of learning historical image features defined by Equation (27) is deactivated, facilitating ablation and analysis of this mechanism. The asterisk-marked results demonstrate that learning historical image features is effective in mapping out constructive hypothetical options.
XI-C Ablation Study on WORLD
The ablation study on WORLD focuses on the hyperparameter $M$. $M$ indicates the number of components in the GMM constructed by World for RPM image features. The value of $M$ determines the maximum diversity of solution distributions that WORLD can generate for DIO, expressed as $M^{N}$. Therefore, ablation experiments centered around the value of $M$ are warranted. The reasoning accuracies of DIO+WORLD on the key sub-tasks of RAVEN, I-RAVEN, and PGM under different $M$ values are recorded in Table IX.
TABLE IX: Accuracies of DIO+WORLD under Different $M$.
Dataset Sub-task Accuracy(%)
MM=512 MM=256 MM=128 MM=32 Solo DIO
RAVEN/
I-RAVEN
$\times3$ Grid 98.8 / 99.5 98.5 / 99.1 97.9 / 98.5 96.5 / 97.9 96.1 / 97.4
O-IG 99.0 / 99.5 98.6 / 99.0 98.0 / 98.7 97.1 / 98.0 96.7 / 97.9
PGM Neutral 99.3 99.0 98.5 98.0 98.1
The number of components directly affects the complexity of the GMM. WORLD proposes an assumption for DIO, stating that the RPM image features encoded by DIO should conform to a GMM. As a highly versatile distribution model, the GMM can theoretically ensure WORLD’s broad applicability. However, as shown in Table IX, when the number of components is relatively small ($M=32$), WORLD constructs a less complex GMM, which negatively impacts DIO’s performance in the PGM task. This paper contends that this phenomenon arises because a 32-component GMM fails to capture the complexity of the underlying distribution of PGM image features. If DIO attempts to fit these more complex features into such a simplified GMM, a performance decline can be anticipated. Therefore, this study establishes a critical prerequisite for WORLD’s applicability: the GMM modeled by WORLD must be able to adapt to the inherent complexity of the target problem.
XII Conclusion
@@
This study has constructed a complete abstract reasoning solution through systematic innovations, with four core innovations forming a progressive technical system that achieves significant progress in both theoretical breakthroughs and empirical performance:
XII-1 DIO
@@
We propose that structuring neural network architectures around the causal chains can lead to enhanced reasoning performance. We also suggest that the learning objectives of end-to-end reasoning models, including DIO, are essentially variational lower bounds of the mutual information between the contexts and the correct options. Consequently, both performance and learning efficacy are directly impacted by the looseness of these lower bounds.
XII-2 Brando
@@
XII-3 WORLD
We constructed a GMM for the distribution of RPM image features, thereby enabling targeted sampling of infinite and semantically rich options. This approach not only assists the DIO model in tightening mutual information but also equips it with the capability to solve open-ended generative RPM problems.
XII-4 DIEGO
We have introduced external metadata to construct supervision signals, precisely correcting semantic biases in the causal chain modeled by DIO. We have established mappings between model reasoning chains and ground-truth causal chains, aligning internal representations with human cognitive frameworks.
Combining these four innovations yields a complete technical closed loop of “Architecture Design $\rightarrow$ Training Optimization $\rightarrow$ Data Augmentation $\rightarrow$ Causal Rectification”.

### v7

@@ X-B Experiment on PGM Problem
Column 1 of Table II summarizes the performance comparison among the four DIO configurations and existing methods on the i.i.d. sub-task “neutral” of the PGM dataset.
TABLE II: Accuracies on PGM Generalization Tasks.
Model Dataset
Neutral Interpolation Extrapolation
@@
CoPINet[18] 56.4 - - - - - - -
SAVIR-T [22] 91.2 - - - - - - -
RS-CNN[26] 82.8 - - - - - - -
MRNet[20] 94.5 68.1 19.2 16.9 30.1 25.9 55.3 38.4
RS-Tran[26] 97.5 77.2 19.2 12.9 24.7 22.2 43.6 28.4
DIO 98.1 82.0 18.8 13.6 26.5 23.3 45.5 29.9
DIO+Brando 99.0 83.2 18.2 14.2 26.7 23.6 46.8 32.8
DIO+WORLD 99.3 85.4 18.9 15.0 27.6 23.8 55.5 39.0
DIO+WORLDG 99.5 87.8 18.9 15.0 27.6 23.8 56.2 39.5
DIO+DIEGO 99.6 93.2 13.8 14.4 27.2 28.8 98.3 98.3
As shown in this column, DIO attains SOTA performance on PGM-neutral, a benchmark on which earlier models achieved relatively low accuracy. Notably, incorporating mutual information lower bound tightening methods (Brando and WORLD) yields substantial improvements for DIO. Moreover, DIO+DIEGO reaches a peak accuracy of 99.6%, challenging the prevailing conclusion that direct metadata supervision necessarily impairs relational reasoning.
Columns 2 onward in Table II record the experimental results on the o.o.d. sub-tasks of PGM. All results in Table II jointly show that mutual-information tightening and causal-chain modeling yield comparable gains for DIO on i.i.d. RPM tasks, while the latter delivers significantly better generalization under o.o.d. conditions. This suggests that, although mutual information as a statistical measure does not inherently encode the causal relationships vital for reasoning, it remains highly effective for i.i.d. tasks; however, when tackling o.o.d. challenges, adopting causal-chain modeling as the core methodology is the more robust choice. We suggest that this occurs because, under i.i.d. conditions, the correlations between contexts and solutions coincide with their causal relationships, allowing mutual-information-driven models to remain effective; in contrast, o.o.d. reasoning tasks expose a divergence between correlation and causality, making causal-chain-driven approaches essential.
@@ X-C Experiment on Generative RAVEN and PGM Problems
We argue that a progressive-pattern-driven discriminative model is a reliable way to verify solutions from generative RPM tasks. Therefore, we use DIO+DIEGO to evaluate the generation quality of DIO+WORLDG. Before running this pipeline, we must report DIEGO’s discriminative performance for progressive patterns under RPM metadata supervision to confirm its reliability as a solution validator. Table III lists DIEGO’s reasoning accuracy for progressive patterns on RAVEN, I-RAVEN, and PGM.
TABLE III: DIEGO’s Reasoning Accuracy for Progressive Patterns on the RAVEN, I-RAVEN, and PGM Datasets.
@@
Table III shows that DIO+DIEGO achieves extremely high accuracy on progressive-pattern reasoning, making it a suitable validator for generative RPM tasks. Consequently, we use DIO+DIEGO to evaluate DIO+WORLDG and report its accuracy on generative RAVEN, I-RAVEN, and PGM-neutral problems in Table IV.
TABLE IV: Generation Accuracies of DIO+WORLDG
@@
The relevant generation results are provided in Figure 11, where, for RAVEN, we display only the most typical $\times3$ Grid and O-IG sub-task outcomes.
Figure 11: Generated RPM images.
@@
Table IV shows that, without extra supervisory signals, DIO+WORLDG has attained competitive results in open-ended generative RPM problems, strongly confirming its significant research value.
X-D Semantic Disentanglement via WORLD Encoding.
Under the WORLD framework, as illustrated in Figure 7, each RPM image $x_{i}$ is encoded into two latent features: $u_{ij}$ and $b_{ij}\sim\mathcal{N}(0,1)$, and the downstream reasoning operates on $u_{ij}+\epsilon_{ij}$. This may give rise to a latent division of labor: $u_{ij}$ encapsulates the core semantics required for reasoning, while the decoder $\mathcal{D}(\cdot)$, trained to maximize the mutual information between $u_{ij}+b_{ij}$ and $x_{i}$, pushes $b_{ij}$ to carry task-irrelevant perturbations. To verify this hypothesis, we compare the reconstructed image $\mathcal{D}(u_{ij}+\epsilon_{ij})$ with the original $x_{i}$ in terms of semantics, revealing how random variations in $b_{ij}$ drive semantic drift. The comparisons are shown in Figure 12.
Figure 12: Verification on Image Semantics.
As can be seen in Figure 12, DIO+WORLDG begins to disentangle perturbing semantics, such as the occlusion relationships among background lines in PGM and the rotation angles of entities in RAVEN, from $u_{ij}$ to some extent.
XI Conclusion
@@
This study has constructed a comprehensive abstract reasoning solution through systematic innovations, with four core innovations forming a progressive technical system that achieves significant advancements in both theoretical breakthroughs and empirical performance:
XI-1 DIO
@@
We propose that structuring networks around causal chains can enhance reasoning performance. We further suggest that the learning objectives of end-to-end reasoning models, including DIO, are essentially variational lower bounds on the mutual information between contexts and correct options. Consequently, however elaborate the network, performance hinges on the tightness of this bound. This points out that the key to machine intelligence lies in mutual-information tightening or causal-chain modeling.
XI-2 Brando
@@
XI-3 WORLD
We construct a GMM for the distribution of RPM image features, enabling targeted sampling of infinite and semantically rich options. This approach not only helps the DIO model tighten mutual information and, to some extent, disentangle perturbing semantics, but also equips it with the capability to solve open-ended generative RPM problems.
XI-4 DIEGO
We introduce a non-statistical constraint that precisely corrects semantic biases in the causal chain modeled by DIO, and we establish mappings between the model’s reasoning chains and the ground-truth causal chains, aligning internal representations with human cognitive frameworks.
This work unites four causal-information innovations that together yield a complete abstract-reasoning stack.